# Shaking Up VLMs: Comparing Transformers and Structured State Space Models for Vision & Language Modeling

### v1

@@ 3.2 Model Architecture
Table 2: Results on visual grounding, and text-oriented, benchmarks. denotes zero-shot performance.
@@ 5.1 Experimental Setup
Similar to previous work (Liu et al., 2024a; Li et al., 2024) we employ a two-step training regime. First, we perform a warmup stage where we train only the connector component on the pretraining dataset. Next, we unfreeze the language model parameters and train on the instruction-tuning dataset. All models are trained using the same data, in the same order, and with identical training hyperparameters (see Appendix B for further details). Unless stated otherwise, we report the evaluation performance without task-specific fine-tuning. The evaluation metrics for each dataset are listed in Table 9.
@@ 5.2 Results
Figure 3: Results of finetuned 1.4B models with increased resolution on VQAv2 (top) and RefCOCOg (bottom). Increasing the resolution in both models results in better performance, however, Pythia benefits significantly more than Mamba in the grounding task.
@@
It is widely known that increasing the image resolution yields benefits in Transformer-based VLMs (Karamcheti et al., 2024; Laurençon et al., 2024b). We explore whether the benefits of higher image resolution translate to Mamba given its strong long sequence modeling capabilities (Gu and Dao, 2023). Figure 3 shows the performance of 1.4B models on VQAv2 and RefCOCOg after finetuning on each task with higher-resolution images. As expected, both models benefit from higher-resolution images, and the differences are more evident in RefCOCOg, possibly due to the granularity of the task. Comparing Pythia-VL and Mamba-VL, both models exhibit a similarly small improvement in VQAv2, but Pythia-VL benefits substantially more than Mamba-VL in RefCOCOg. This provides further evidence regarding the limitations of Mamba on grounding tasks, on which we further elaborate in Section 5.3.
@@
For completeness, we provide a comparison against state-of-the-art 3B and 7B parameter models (Table 3). We observe that our largest models are competitive even against the largest VLM models. Importantly, we want to note that our LLM has not been instruction-tuned beforehand, which could have a major impact, particularly in multimodal language modeling tasks (Laurençon et al., 2024b). Furthermore, it is hard to draw definite conclusions between different models as they have been trained using different datasets and training regimes.
@@ 6 Conclusion
It has been increasingly transparent that the cost of training large-scale models, including VLMs, raises compute barriers Strubell et al. (2019); Thompson et al. (2020); Luccioni et al. (2024). While patch representations have become the standard approach for encoding images, these representations substantially increase the context window and, consequently, the computational cost of training. To improve efficiency, we have employed sequence packing during training, which results to fewer padding tokens within the batch. Additionally, more sophisticated V&L connectors that downsmple the visual sequence (Alayrac et al., 2022; Dai et al., 2024; Laurençon et al., 2024a) can, in principle, accelerate training and inference. We leave comparisons of more efficient V&L connectors in combination with SSMs as future work.
@@
A widely acknowledged limitation for LLMs and VLMs is the factuality of the generated content Ji et al. (2023). The impact of this property can vary depending on the downstream task (e.g., answering a question accurately versus creating novel images with text prompts). In this work, we use POPE (Li et al., 2023), a benchmark specifically designed to evaluate object hallucinations in VLMs. However, further investigation is needed to evaluate model hallucinations and improve the reliability of VLMs.

### v2

@@ 3.2 Model Architecture
Table 2: Results on visual grounding, and text-oriented, benchmarks. $\dagger$ denotes a task not in the training mixture.
@@ 5.1 Experimental Setup
Similar to previous work (Liu et al., 2024a; Li et al., 2024) we employ a two-step training regime. First, we perform a warmup stage where we train only the connector component on the pretraining dataset. Next, we unfreeze the language model parameters and train on the instruction-tuning dataset. All models are trained using the same data, in the same order, and with identical training hyperparameters (see Appendix B for further details). Unless stated otherwise, we report the evaluation performance without task-specific fine-tuning.
@@ 5.2 Results
Figure 3: Results of finetuned 1.4B models with increased resolution on VQAv2 (top), RefCOCOg (middle), and TextVQA (bottom). Increasing the resolution to 480$\times$480 pixels results better performance for both models, however, Pythia benefits significantly more than Mamba in the grounding task.
@@
It is widely known that increasing the image resolution yields benefits in Transformer-based VLMs (Karamcheti et al., 2024; Laurençon et al., 2024b). We explore whether the benefits of higher image resolution translate to Mamba given its strong long sequence modeling capabilities (Gu and Dao, 2023). Figure 3 shows the performance of 1.4B models on VQAv2, RefCOCOg, and TextVQA after finetuning on each task with higher-resolution images. As expected, both models benefit from higher-resolution images, and the differences are more evident in RefCOCOg, possibly due to the granularity of the task. Comparing Pythia-VL and Mamba-VL, both models exhibit a similarly small improvement in VQAv2 and TextVQA, but Pythia-VL benefits substantially more than Mamba-VL in RefCOCOg. This provides further evidence regarding the limitations of Mamba on grounding tasks, on which we further elaborate in Section 5.3.
@@
For completeness, we provide a comparison against state-of-the-art 3B and 7B parameter models (Table 3). We observe that our largest models are competitive even against the largest VLM models. Our model performs on par with other Mamba-based VLMs (Cobra and VL-Mamba) with a small advantage in general VQA benchmarks (VQA, GQA). Importantly, we want to note that our base LLMs have not been instruction-tuned, which could have a major impact, particularly in multimodal language modeling tasks (Laurençon et al., 2024b). Furthermore, it is hard to draw definite conclusions between different models as they have been trained using different datasets and training regimes.
@@ 6 Conclusion
It has been increasingly transparent that the cost of training large-scale models, including VLMs, raises compute barriers Strubell et al. (2019); Thompson et al. (2020); Luccioni et al. (2024). While patch representations have become the standard approach for encoding images, these representations substantially increase the context window and, consequently, the computational cost of training. To improve efficiency, we have employed sequence packing during training, which results to fewer padding tokens within the batch. Additionally, more sophisticated V&L connectors that downsample the visual sequence (Alayrac et al., 2022; Dai et al., 2024; Laurençon et al., 2024a) can, in principle, accelerate training and inference. We leave comparisons of more efficient V&L connectors in combination with SSMs as future work.
@@
A widely acknowledged limitation for LLMs and VLMs is the factuality of the generated content Ji et al. (2023). The impact of this property can vary depending on the downstream task (e.g., answering a question accurately versus creating novel images with text prompts). Furthermore, prior work (Pantazopoulos et al., 2024), has shown that the visual instruction tuning stage imposes a forgetting effect on the safety guardrails of the backbone LLM leading to more vulnerable VLMs. In this work, we use POPE (Li et al., 2023), a benchmark specifically designed to evaluate object hallucinations in VLMs. However, further investigation is needed to evaluate model hallucinations and improve the reliability of VLMs.
Acknowledgements
We would like to thank the reviewers for their valuable feedback during the ARR process. Additionally, this work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh. Finally, the authors acknowledge the use of the HWU high-performance computing facility (DMOG) and associated support services in the completion of this work.